# Submodular Maximization over a Matroid $k$-Intersection: Multiplicative Improvement over Greedy

### v1

@@ 4 Analysis of the Main Algorithm
In this section, we bound the approximation guarantee of Algorithm 1 when $f\colon 2^{E}\to{\mathbb{R}}$ is a general submodular function with $f(\emptyset)\geq 0$. We then use this bound to prove Theorem $1.$ Fix a matroid $k$-parity instance $(E,{\mathcal{I}})$. In the following, we consider an execution of Algorithm 1 on this matroid $k$-parity and the above function $f$. Recall that $A_{\leq L}$ is the output of this algorithm, and thus, our objective is to lower bound $f(A_{\leq L})$ in terms of $f(O)$, where $O$ is an optimal solution for our given instance (i.e., a set in $\mathcal{I}$} maximizing $f$). Without loss of generality, we may restrict our analysis to optimal solutions that satisfy the following property.
@@ 4.1 Partitioning the Elements of $O$
This section describes a partition of $O$ into disjoint sets $O_{1},O_{2},\dotsc,O_{L}$. Analogously to the sets $A_{\leq i}$, for every integer $0\leq i\leq L$, we use $O_{\leq i}$ to denote the union $\cup_{j=1}^{i}O_{j}$. Intuitively, each set $O_{i}$ represents elements from $O$ whose value can be bounded in terms of the elements selected as $A_{i}$ in iteration $i$ of Algorithm 1, and thus, need not be considered when later iterations of this algorithm are analyzed. In line with this intuition, we prove below that the elements of $O\setminus O_{\leq i}$, whose value needs to be bounded in terms of elements selected by later iterations, can all be feasibly added to the solution $A_{\leq i}$ that Algorithm 1 has after iteration ii.
@@
where $\{N_{o}^{(i)}\subseteq A_{\leq i}\mid o\in A_{i-1}\cup(O\setminus O_{\leq{i-1}})\}$ is the collection of sets obtained by applying Lemma 2.2 with the sets $A_{\leq i}$ and $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ as the sets AA and $B$ of the lemma, respectively. Recall that applying Lemma 2.2 with these sets requires them to be feasible. The feasibility of $A_{\leq i}$ is clear. The feasibility of $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ follows from the next observation.
@@
We prove the observation by induction on ii. First, note that the observation is trivial for $i=0$ because $A_{\leq 0}\cup(O\setminus O_{\leq 0})=O$ and $A_{\leq 0}=\emptyset$. Assume now that the observation holds for every integer $0\leq i^{\prime}<i$, and let us prove it for ii.
@@ 4.2 Defining Weights for the Charging Argument
We now define the weights we use to quantify the values of elements in $O$ and $A_{\leq L}$. An element’s weight is based on the incremental value it adds to a given set, and in the linear case it corresponds exactly to the weight of this element in the definition the objective function $f$ (as long as this weight is non-negative). In order to define the incremental values, we order the elements of $A_{\leq L}$ as $a_{1},a_{2},\dotsc,a_{|A_{\leq L}|}$, in the order in which they were last added by Algorithm 1 to any of the sets $A_{1},A_{2},\dotsc,A_{L}$. Applying $(\theta,\varepsilon)$-improvements of the third kind involves adding 22 elements to $A_{i}$ at the same time, which creates an ambiguity. To resolve the ambiguity, recall that in such improvements, Definition 3.1 specifies that there must be a way to assign the two added elements the labels $x_{1}$ and $x_{2}$ so that $f(x_{1}\mid A)\geq\theta$ and $f(x_{2}\mid A+x_{1})\geq\theta$, where AA is the set improved by the $(\theta,\varepsilon)$-improvement. In our ordering of $A_{\leq L}$, we suppose that the element corresponding to $x_{1}$ is added to $A_{i}$ immediately before the element corresponding to $x_{2}$. [^4] If both ways to label the added elements obey the above inequalities, then we arbitrary select one of them. Given our ordering $a_{1},\dotsc,a_{|A_{\leq L}|}$, we define the weight of each element $a_{j}\in A_{\leq L}$ as
@@ 4.3 Classifying the Elements of $O$ for Charging
for any $j^{\prime}<j$. Intuitively, these inequalities show that any such $o_{j}$ has marginal contribution large enough to be part of an $(m_{i},\varepsilon)$-exchange at the end of the iteration ii. Using these observations, we now prove each claim of the lemma in turn.
@@
Fix any $a\in A_{\leq L}$ and let $O^{\prime}=\{o\in O_{\leq L}\mid a\in N_{o}\}$. By the same argument as in the previous claim, since the sets $A_{i}$ are disjoint and $N_{o}^{(i)}\subseteq A_{i}$ for all $i\in[L]$, there must be some single $i\in[L]$ such that $a\in A_{i}$ and $N_{o}=N_{o}^{(i)}$ for all $o\in O^{\prime}$. For this $i$, the sets $\{N_{o}^{(i)}\mid o\in A_{\leq i-1}\cup(O\setminus O_{\leq i-1})\}$ were constructed to satisfying the properties of Lemma $2.$ By the fourth such property, $a$ appears in at most $k$ of them, and hence, $|O^{\prime}|\leq k$. ∎
@@ 4.5 The Charging Scheme
Next, we would like to bound $u(o)$ for the elements in $O_{\leq L}\setminus O^{(s)}$. Unlike the bounds in the previous two lemmata, this bound depends on the random shift $\tau$ selected by the algorithm. To formulate this bound, let us define, for each $o\in O$, $m^{(o)}\triangleq\min\{m_{i}\mid i\geq 0,m_{i}\geq u(o)\}$ to be the value of the smallest threshold $m_{i}$ that is at least $u(o)$ (observe that $m^{(o)}$ is well-defined since $u(o)>0$ by Observation 4.4 and $u(o)\leq W\leq m_{0}$ by submodularity). Additionally, we define the ratio $r_{o}\triangleq m^{(o)}/u(o)$. Notice that $r_{o}$ is always in the range $[1,2)$ since $u(o)\leq m^{(o)}$ and $m_{i-1}/m_{i}=2$ for all ii. Given this notation, we can now state the promised bound as Lemma 4.8. This lemma introduces a parameter $d\geq 1$, which is later used to control the balance between the bound of this lemma and the aforementioned bound from Lemma 4.5.
@@
Assuming $f$ is a submodular function such that $f(\emptyset)\geq 0$, the output set $A_{\leq L}$ of Algorithm 1 obeys, for every $d\geq d^{\prime}\geq 1$ and any strictly down-monotone $O\in{\mathcal{I}}$,
[…]
@@
Combining $d^{\prime}$ times the bounds from Lemmata 4.6 and 4.7 with the bound from Lemma 4.8, we obtain
@@
$\leq\sum_{\mathclap{o\in O_{\leq L}}}w(N_{o})+((1+\varepsilon)d^{\prime}-1)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq\sum_{\mathclap{a\in A_{\leq L}}}k\cdot w(a)+((1+\varepsilon)d^{\prime}-1)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq(k+(1+\varepsilon)d^{\prime}-1)\cdot f(A_{\leq L}\mid\emptyset)+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]\kern 5.0pt,$
where the second inequality follows from the condition $d\geq d^{\prime}\geq 1$ and the inequality $u(o)-{\bar{w}}(o)\geq 0$ (which holds by Observation 4.4); the third inequality follows from the fact that each $a\in A_{\leq L}$ appears in at most $k$ sets $N_{o}$ (claim 3 of Lemma 4.3) and obeys $w(a)\geq 0$ (claim 1 of Lemma 4.2); and the final inequality follows from the inequality $\sum_{o\in O^{(s)}}w(N_{o})\leq f(A_{\leq L}\mid\emptyset)$ proved by Lemma 4.6.
@@
[…]
$\leq d^{\prime}\cdot f(\emptyset)+(k+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(A_{\leq L}\mid\emptyset)]$
@@
When $f$ is a linear function, the approximation guarantee of Algorithm 1 is at most $k\ln 2+1-\ln 2+\varepsilon$.
@@
[…]
Let $d^{\prime}=\ln^{-1}2$. Below, we prove that ${\mathbb{E}}[r_{o}]=d^{\prime}$ for every $o\in O$, which implies that the last term on the rightmost side of the above inequality is 0. Rearranging the above inequality then yields
${\mathbb{E}}[f(A_{\leq L})]\geq\frac{d^{\prime}}{k+(1+\varepsilon)d^{\prime}-1}\cdot f(O)=\frac{\ln^{-1}2}{k+(1+\varepsilon)\ln^{-1}2-1}\cdot f(O)=\frac{f(O)}{k\ln 2+1-\ln 2+\varepsilon}\kern 5.0pt.$
@@
$(k+(1+\varepsilon)d^{\prime}+d-1)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)+d\cdot{\mathbb{E}}[f(A_{\leq L}\mid O)]\kern 5.0pt,$
@@
One can verify that $d\geq d^{\prime}\geq 1$ for every $d\geq 1$. Proposition 4.9 and Lemma 4.5 thus give together
[…]
@@
$(k+(1+\varepsilon)d^{\prime}+d-1)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)\kern 5.0pt,$
@@
$\frac{k+d+(1+\varepsilon)\cdot\Big(\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}\Big)-1}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}}=1+\varepsilon+\frac{k+d-1}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}}\\ =1+\varepsilon+\frac{k+d-1}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}}\leq 1+\varepsilon+\frac{(k+d-1)\cdot\big(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2}\big)}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the inequality uses the fact that for every two positive values $a,b$ obeying $2b\leq a$, we have $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Selecting $d=\sqrt{k}$, and recalling that $\varepsilon<1$, makes the above expression equal to $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})$, which proves the corollary since there must exist some optimal solution $O$ that is also strictly down-monotone. ∎
@@ 5 Algorithm for Non-monotone Submodular Functions
In this section, we present and analyze Algorithm 2, which we use to prove Theorem $1.$ This algorithm receives as input a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$, a matroid $k$-parity constraint $(E,{\mathcal{I}})$, and a positive integer parameter $\ell$. Algorithm 2 runs for $\ell$ iterations. In each iteration $i$, Line 2 considers the restriction of our problem to some ground set $E_{i-1}\subseteq E$. Note that the restricted $f_{i}$ inherits the submodularity of $f$ and the resulting constraint $(E_{i-1},2^{E_{i-1}}\cap{\mathcal{I}})$ is also a matroid $k$-parity constraint. In each iteration, we apply Algrithm 1 to obtain a feasible set $B_{i}\subseteq E_{i-1}$. Next, in Line 2 the algorithm invokes the Double Greedy algorithm of Buchbinder et al. [5]. This algorithm takes a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ and a set $S\subseteq E$, and returns a set $T\subseteq S$ such that ${\mathbb{E}}[f(T)]\geq\frac{1}{2}\cdot\max_{S^{\prime}\subseteq S}f(S^{\prime})$. We denote by $B^{\prime}_{i}$ the output of Double Greedy applied to our set $B_{i}$. Algorithm 2 then returns as its output the best set constructed amongst all the sets $B_{i}$ and $B^{\prime}_{i}$ in all $\ell$ iterations.
@@
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot{\mathbb{E}}[f(O\cap E_{i-1})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
@@
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot f(O^{\prime})+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O^{\prime})]$ (5)
@@
$(k+d+(2i-1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]\geq d^{\prime}\cdot{\mathbb{E}}[f(O)]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
@@
$(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]\geq{}$ $(k+d+(1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B_{i})]$
@@
$\ell\cdot(k+d+(\ell+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]={}$ $\sum_{i=1}^{\ell}(k+d+(2i-1+\varepsilon)d^{\prime}-1)\cdot{\mathbb{E}}[f(B^{*})]$
@@
where the second inequality follows from the submodularity of $f$, and the last inequality follows from $f$’s non-negativity. Recall that there must be some strictly down-montone set $O$ that is also an optimal solution. Thus, the approximation ratio of Algorithm 2 is at most
$\frac{\ell\cdot(k+d+(\ell+\varepsilon)d^{\prime}-1)}{d^{\prime}\ell-d}={}$ $\ell+\varepsilon+\frac{\ell\cdot(k+2d-1)+\varepsilon d}{d^{\prime}\ell-d}$
$={}$ $\ell+\varepsilon+\frac{k+2d-1+\varepsilon d/\ell}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}-d/\ell}=\ell+\varepsilon+\frac{k+2d-1+\varepsilon d/\ell}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}-d/\ell}$
$\leq{}$ $\ell+\varepsilon+\frac{(k+2d-1+\varepsilon d/\ell)\cdot\big(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2}+\frac{2d}{\ell}\cdot\frac{2\ln 2}{1+\ln 2}\big)}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the inequality holds for $\ell\geq 3d$ because for every two positive values $a,b$ obeying $2b\leq a$, it holds that $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Choosing now $\ell=\lceil 3k^{2/3}\rceil$ and $d=k^{1/3}$, and recalling that $\varepsilon$ is assigned a constant value from $(0,1)$, we get that the approximation ratio of Algorithm 2 is at most $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})$.

### v2

@@ 4 Analysis of the Main Algorithm
In this section, we bound the approximation guarantee of Algorithm 1 when $f\colon 2^{E}\to{\mathbb{R}}$ is a general submodular function with $f(\emptyset)\geq 0$. We then use this bound to prove Theorem 1.1. Fix a matroid $k$-parity instance $(E,{\mathcal{I}})$. In the following, we consider an execution of Algorithm 1 on this matroid $k$-parity and the above function $f$. Recall that $A_{\leq L}$ is the output of this algorithm, and thus, our objective is to lower bound $f(A_{\leq L})$ in terms of $f(O)$, where $O$ is an optimal solution for our given instance (i.e., a set in $\mathcal{I}$} maximizing $f$). Without loss of generality, we may restrict our analysis to optimal solutions that satisfy the following property.
@@ 4.1 Partitioning the Elements of $O$
This section describes a partition of $O$ into disjoint sets $O_{1},O_{2},\dotsc,O_{L}$. Analogously to the sets $A_{\leq i}$, for every integer $0\leq i\leq L$, we use $O_{\leq i}$ to denote the union $\cup_{j=1}^{i}O_{j}$. Intuitively, each set $O_{i}$ represents elements from $O$ whose value can be bounded in terms of the elements selected as $A_{i}$ in iteration $i$ of Algorithm 1, and thus, need not be considered when later iterations of this algorithm are analyzed. In line with this intuition, we prove below that the elements of $O\setminus O_{\leq i}$, whose value needs to be bounded in terms of elements selected by later iterations, can all be feasibly added to the solution $A_{\leq i}$ that Algorithm 1 has after iteration $i$.
@@
where $\{N_{o}^{(i)}\subseteq A_{\leq i}\mid o\in A_{i-1}\cup(O\setminus O_{\leq{i-1}})\}$ is the collection of sets obtained by applying Lemma 2.2 with the sets $A_{\leq i}$ and $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ as the sets $A$ and $B$ of the lemma, respectively. Recall that applying Lemma 2.2 with these sets requires them to be feasible. The feasibility of $A_{\leq i}$ is clear. The feasibility of $A_{\leq i-1}\cup(O\setminus O_{\leq i-1})$ follows from the next observation.
@@
We prove the observation by induction on $i$. First, note that the observation is trivial for $i=0$ because $A_{\leq 0}\cup(O\setminus O_{\leq 0})=O$ and $A_{\leq 0}=\emptyset$. Assume now that the observation holds for every integer $0\leq i^{\prime}<i$, and let us prove it for $i$.
@@ 4.2 Defining Weights for the Charging Argument
We now define the weights we use to quantify the values of elements in $O$ and $A_{\leq L}$. An element’s weight is based on the incremental value it adds to a given set, and in the linear case it corresponds exactly to the weight of this element in the definition the objective function $f$ (as long as this weight is non-negative). In order to define the incremental values, we order the elements of $A_{\leq L}$ as $a_{1},a_{2},\dotsc,a_{|A_{\leq L}|}$, in the order in which they were last added by Algorithm 1 to any of the sets $A_{1},A_{2},\dotsc,A_{L}$. Applying $(\theta,\varepsilon)$-improvements of the third kind involves adding $2$ elements to $A_{i}$ at the same time, which creates an ambiguity. To resolve the ambiguity, recall that in such improvements, Definition 3.1 specifies that there must be a way to assign the two added elements the labels $x_{1}$ and $x_{2}$ so that $f(x_{1}\mid A)\geq\theta$ and $f(x_{2}\mid A+x_{1})\geq\theta$, where $A$ is the set improved by the $(\theta,\varepsilon)$-improvement. In our ordering of $A_{\leq L}$, we suppose that the element corresponding to $x_{1}$ is added to $A_{i}$ immediately before the element corresponding to $x_{2}$. [^4] If both ways to label the added elements obey the above inequalities, then we arbitrary select one of them. Given our ordering $a_{1},\dotsc,a_{|A_{\leq L}|}$, we define the weight of each element $a_{j}\in A_{\leq L}$ as
@@ 4.3 Classifying the Elements of $O$ for Charging
for any $j^{\prime}<j$. Intuitively, these inequalities show that any such $o_{j}$ has marginal contribution large enough to be part of an $(m_{i},\varepsilon)$-exchange at the end of the iteration $i$. Using these observations, we now prove each claim of the lemma in turn.
@@
Fix any $a\in A_{\leq L}$ and let $O^{\prime}=\{o\in O_{\leq L}\mid a\in N_{o}\}$. By the same argument as in the previous claim, since the sets $A_{i}$ are disjoint and $N_{o}^{(i)}\subseteq A_{i}$ for all $i\in[L]$, there must be some single $i\in[L]$ such that $a\in A_{i}$ and $N_{o}=N_{o}^{(i)}$ for all $o\in O^{\prime}$. For this $i$, the sets $\{N_{o}^{(i)}\mid o\in A_{\leq i-1}\cup(O\setminus O_{\leq i-1})\}$ were constructed to satisfying the properties of Lemma 2.2. By the fourth such property, $a$ appears in at most $k$ of them, and hence, $|O^{\prime}|\leq k$. ∎
@@ 4.5 The Charging Scheme
Next, we would like to bound $u(o)$ for the elements in $O_{\leq L}\setminus O^{(s)}$. Unlike the bounds in the previous two lemmata, this bound depends on the random shift $\tau$ selected by the algorithm. To formulate this bound, let us define, for each $o\in O$, $m^{(o)}\triangleq\min\{m_{i}\mid i\geq 0,m_{i}\geq u(o)\}$ to be the value of the smallest threshold $m_{i}$ that is at least $u(o)$ (observe that $m^{(o)}$ is well-defined since $u(o)>0$ by Observation 4.4 and $u(o)\leq W\leq m_{0}$ by submodularity). Additionally, we define the ratio $r_{o}\triangleq m^{(o)}/u(o)$. Notice that $r_{o}$ is always in the range $[1,2)$ since $u(o)\leq m^{(o)}$ and $m_{i-1}/m_{i}=2$ for all $i$. Given this notation, we can now state the promised bound as Lemma 4.8. This lemma introduces a parameter $d\geq 1$, which is later used to control the balance between the bound of this lemma and the aforementioned bound from Lemma 4.5.
@@
Assuming $f$ is a submodular function such that $f(\emptyset)\geq 0$, the output set $A_{\leq L}$ of Algorithm 1 obeys, for every two values $d\geq 2$ and $d^{\prime}\leq k+1$ and any strictly down-monotone $O\in{\mathcal{I}}$,
$(k+1+2\varepsilon)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)-d\cdot\sum_{o\in O}{\mathbb{E}}[u(o)-{\bar{w}}(o)]-\sum_{o\in O}[d^{\prime}-\operatorname*{\mathbb{E}}[\rho_{o,d}]]\cdot u(o)\kern 5.0pt,$
@@
Combining $\rho_{o,d}$ times the bounds from Lemmata 4.6 and 4.7 with the bound from Lemma 4.8, we obtain
@@
$\leq\sum_{\mathclap{o\in O_{\leq L}}}w(N_{o})+(1+2\varepsilon)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq\sum_{\mathclap{a\in A_{\leq L}}}k\cdot w(a)+(1+2\varepsilon)\sum_{\mathclap{o\in O^{(s)}}}w(N_{o})+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]$
$\leq(k+1+2\varepsilon)\cdot f(A_{\leq L}\mid\emptyset)+d\cdot\sum_{o\in O}[u(o)-{\bar{w}}(o)]\kern 5.0pt,$
where the first inequality uses the facts that $\rho_{o,d}\leq r_{o}\leq 2$ and $u(o)>0$ (by Observation 4.4); the third inequality follows from $d\geq 2$ and the inequality $u(o)-{\bar{w}}(o)\geq 0$ (which holds by Observation 4.4); the fourth inequality follows from the fact that each $a\in A_{\leq L}$ appears in at most $k$ sets $N_{o}$ (claim 3 of Lemma 4.3) and obeys $w(a)\geq 0$ (claim 1 of Lemma 4.2); and the final inequality follows from the inequality $\sum_{o\in O^{(s)}}w(N_{o})\leq f(A_{\leq L}\mid\emptyset)$ proved by Lemma 4.6.
@@
$d^{\prime}\cdot f(O)$ $={d^{\prime}\cdot f(\emptyset)+{\mathbb{E}}\bigg[\sum_{\mathclap{o\in O}}\rho_{o,d}\cdot u(o)\bigg]+{\mathbb{E}}\bigg[\sum_{o\in O}(d^{\prime}-\rho_{o,d})\cdot u(o)\bigg]}$
@@
When $f$ is a linear function, the approximation guarantee of Algorithm 1 is at most $(k+1)\ln 2+O(\varepsilon)$.
@@
$(k+1+2\varepsilon)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)-\sum_{o\in O}[d^{\prime}-{\mathbb{E}}[r_{o}]]\cdot u(o)\kern 5.0pt.$
Let $d^{\prime}=\ln^{-1}2$ (notice that $d^{\prime}\leq 2\leq k+1$). Below, we prove that ${\mathbb{E}}[r_{o}]=d^{\prime}$ for every $o\in O$, which implies that the last term on the rightmost side of the above inequality is 0. Rearranging the above inequality then yields
${\mathbb{E}}[f(A_{\leq L})]\geq\frac{d^{\prime}}{k+1+2\varepsilon}\cdot f(O)=\frac{\ln^{-1}2}{k+1+2\varepsilon}\cdot f(O)=\frac{f(O)}{(k+1)\ln 2+O(\varepsilon)}\kern 5.0pt.$
@@
$(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)+d\cdot{\mathbb{E}}[f(A_{\leq L}\mid O)]\kern 5.0pt,$
@@
Note that $d^{\prime}=\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}\leq\frac{1}{2\ln 2}+\frac{3d/2}{2d}<1+3/4<2$. Thus, Proposition 4.9 and Lemma 4.5 give together
$(k+1+2\varepsilon)\cdot{}$ ${\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)-d\cdot{\mathbb{E}}\Bigl[\sum_{\mathclap{o\in O}}u(o)-{\bar{w}}(o)\Bigr]-\sum_{o\in O}[d^{\prime}-{\mathbb{E}}[\rho_{o,d}]]\cdot u(o)$
$\geq d^{\prime}\cdot f(O)-d\cdot{\mathbb{E}}[f(A_{\leq L})-f(A_{\leq L}\mid O)]-\sum_{o\in O}[d^{\prime}-{\mathbb{E}}[\rho_{o,d}]]\cdot u(o)\kern 5.0pt.$
@@
$(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(A_{\leq L})]\geq d^{\prime}\cdot f(O)\kern 5.0pt,$
@@
$\frac{k+d+1+2\varepsilon}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}}=\frac{k+d+1+2\varepsilon}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}}\leq\frac{(k+d+1+2\varepsilon)(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2})}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the inequality uses the fact that for every two positive values $a,b$ obeying $2b\leq a$, we have $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Selecting $d=2\sqrt{k}$ (notice that $d\geq 2$) makes the above expression equal to $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})$, which proves the corollary since there must exist some optimal solution $O$ that is also strictly down-monotone. ∎
@@ 5 Algorithm for Non-monotone Submodular Functions
In this section, we present and analyze Algorithm 2, which we use to prove Theorem 1.1. This algorithm receives as input a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$, a matroid $k$-parity constraint $(E,{\mathcal{I}})$, and a positive integer parameter $\ell$. Algorithm 2 runs for $\ell$ iterations. In each iteration $i$, Line 2 considers the restriction of our problem to some ground set $E_{i-1}\subseteq E$. Note that the restricted $f_{i}$ inherits the submodularity of $f$ and the resulting constraint $(E_{i-1},2^{E_{i-1}}\cap{\mathcal{I}})$ is also a matroid $k$-parity constraint. In each iteration, we apply Algrithm 1 to obtain a feasible set $B_{i}\subseteq E_{i-1}$. Next, in Line 2 the algorithm invokes the Double Greedy algorithm of Buchbinder et al. [5]. This algorithm takes a non-negative submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ and a set $S\subseteq E$, and returns a set $T\subseteq S$ such that ${\mathbb{E}}[f(T)]\geq\frac{1}{2}\cdot\max_{S^{\prime}\subseteq S}f(S^{\prime})$. We denote by $B^{\prime}_{i}$ the output of Double Greedy applied to our set $B_{i}$. Algorithm 2 then returns as its output the best set constructed amongst all the sets $B_{i}$ and $B^{\prime}_{i}$ in all $\ell$ iterations.
@@
$(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot{\mathbb{E}}[f(O\cap E_{i-1})]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
@@
$(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(B_{i})]\geq{}$ $d^{\prime}\cdot f(O^{\prime})+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O^{\prime})]$ (5)
@@
$(k+d+2d^{\prime}(i-1)+1+2\varepsilon)\cdot{\mathbb{E}}[f(B^{*})]\geq d^{\prime}\cdot{\mathbb{E}}[f(O)]+d\cdot\operatorname*{\mathbb{E}}[f(B_{i}\mid O)]\kern 5.0pt,$
@@
$(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(B^{*})]\geq{}$ $(k+d+1+2\varepsilon)\cdot{\mathbb{E}}[f(B_{i})]$
@@
$\ell\cdot(k+d+d^{\prime}(\ell-1)+1+2\varepsilon)\cdot{\mathbb{E}}[f(B^{*})]={}$ $\sum_{i=1}^{\ell}(k+d+2d^{\prime}(i-1)+1+2\varepsilon)\cdot{\mathbb{E}}[f(B^{*})]$
@@
where the second inequality follows from the submodularity of $f$, and the last inequality follows from $f$’s non-negativity. Recall that there must be some strictly down-montone set $O$ that is also an optimal solution. Thus, when $\ell\geq 2d$, the approximation ratio of Algorithm 2 is at most
$\frac{\ell\cdot(k+d+d^{\prime}(\ell-1)+1+2\varepsilon)}{d^{\prime}\ell-d}\leq{}$ $\ell+\frac{\ell\cdot(k+2d+1+2\varepsilon)}{d^{\prime}\ell-d}$
$={}$ $\ell+\frac{k+2d+1+2\varepsilon}{\frac{1-1/d}{2\ln 2}+\frac{d+1}{2d}-d/\ell}=\ell+\frac{k+2d+1+2\varepsilon}{\frac{1+\ln 2}{2\ln 2}-d^{-1}\cdot\frac{1-\ln 2}{2\ln 2}-d/\ell}$
$\leq{}$ $\ell+\frac{(k+2d+1+2\varepsilon)\cdot\big(1+2d^{-1}\cdot\frac{1-\ln 2}{1+\ln 2}+\frac{2d}{\ell}\cdot\frac{2\ln 2}{1+\ln 2}\big)}{\frac{1+\ln 2}{2\ln 2}}\kern 5.0pt,$
where the second inequality holds because for every two positive values $a,b$ obeying $2b\leq a$, it holds that $\frac{1}{a-b}\leq\frac{1+2b/a}{a}$. Choosing now $\ell=\lceil 4k^{2/3}\rceil$ and $d=2k^{1/3}$, and recalling that $\varepsilon$ is assigned a constant value from $(0,1)$, we get that the approximation ratio of Algorithm 2 is at most $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})$.